A box contains 9 balls numbered from 3 to 11. If 4 random balls are selected (without replacement), what is the probability that the sum of the numbers is even? (explain please)

Thanks for the help
Answer #1

Total number pf ways of selecting 4 balls out of 9 is C(9,4) = 126

Total ways the sum is even is

1. All numbers are even (4,6,8,10) = 1 way

2.

Two numbers are even and two are odd = C(5,2)*C(4,2) = 10*6 = 60 ways

3.

All four number are odd = C(5,4) = 1 way

P(sum of numbers is even) = 62/126 = 0.4921